OpenStudy (anonymous):

Show a graph representing the height of a football t seconds after it was punted. How many seconds did it take the football to land on the ground? Is it A. 4.5 B. 5.5 C. 5 D. 6

OpenStudy (anonymous):

um, You need something to make the graph

OpenStudy (anonymous):

like a function....

OpenStudy (anonymous):

This is the graph it gives

OpenStudy (anonymous):

ok, one sec

OpenStudy (anonymous):

C

OpenStudy (anonymous):

look on the x coordinate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see where it first hits teh x coordinate then the last time it hits it, the time between that is the time

OpenStudy (anonymous):

I have one more I have been stuck on forever can u help me with that one too??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which equation has this graph? A. y = x2 - 6x + 5 B. y = -x2 - 6x + 5 C. y = -x2 - 6x - 5 D. y = x2 + 6x - 5

OpenStudy (anonymous):

are you sure you wrote those right? cause according to me, ( i am using a graphing calc) none fit that graph...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are those supposed to be exponents next to it?

OpenStudy (anonymous):

No it just says Which equation has this graph?

OpenStudy (anonymous):

well, I am graphing them over and over with a calculator and it is not showing that graph for any of them,...

OpenStudy (anonymous):

are you Sure it is not like " y = x^2 - 6x + 5" instead of " y = x2 - 6x + 5"

OpenStudy (anonymous):

Yeah it is x^2 sorry

OpenStudy (anonymous):

i see, ok, let me work it out like that now

OpenStudy (anonymous):

could you please retype them

OpenStudy (anonymous):

Which equation has this graph? A. y = x^2 - 6x + 5 B. y = -x^2 - 6x + 5 C. y = -x^2 - 6x - 5 D. y = x^2 + 6x - 5

OpenStudy (anonymous):

i am not sure, none are working for me...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

its all good up in my hood